OpenStudy (anonymous):

1 + 1 = 3 + 2 = 5? How is it possible?

hartnn (hartnn):

is that , in any way, a serious math question ?

OpenStudy (anonymous):

no no it is not

OpenStudy (anonymous):

Yes, because I want to see if y'all know why!!

OpenStudy (anonymous):

It is not possible

OpenStudy (anonymous):

Its not true....

OpenStudy (anonymous):

cause 1+1=2+2=4

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

1+1 = 5 only when we assume 5=2 Hope it Helps :D

ganeshie8 (ganeshie8):

i dont get it :(

OpenStudy (anonymous):

@Jonathan1997 u made this up

OpenStudy (anonymous):

@hartnn is just on his periond and is pissy

hartnn (hartnn):

oh wow!

hartnn (hartnn):

how did u know ?

OpenStudy (anonymous):

Well, you delete my post before I can even do anything!!

OpenStudy (anonymous):

@Jonathan1997 he can suspend u